OpenStudy (anonymous):

Which of the following equations have the same solution? Explain each step in solving each equation using complete sentences. I.2x + 4 = 3x - 11 II. 3x + 3 = 51 III. 2x+1/3=x+6/2

OpenStudy (anonymous):

andy ideas @texaschic101 ?

OpenStudy (texaschic101):

Equation 1 2x + 4 = 3x - 11 subtract 3x from both sides 2x - 3x + 4 = -11 subtract 4 from both sides 2x - 3x = -11 - 4 combine like terms -x = - 15 multiply by -1 to make x positive x = 15 Equation 2 3x + 3 = 51 subtract 3 from both sides 3x = 51 - 3 3x = 48 divide by 3 x = 16 Equation 3 2x + 1/3 = x + 6/2 subtract x from both sides 2x - x + 1/3 = 6/2 subtract 1/3 from both sides 2x - x = 6/2 - 1/3 common denominator is 6 x = 18/6 - 2/6 x = 16/6 reduces to 8/3 I am not getting the same solution on any of them

OpenStudy (texaschic101):

Equation 3 (2x + 1)/3 = (x + 6)/2 cross multiply 3(x + 6) = 2(2x + 1) distribute through the parenthesis 3x + 18 = 4x + 2 subtract 4x from both sides 3x - 4x + 18 = 2 subtract 18 from both sides 3x - 4x = 2 - 18 -x = -16 multiply by -1 to make x positive x = 16 I found it now...the 2 equations that have the same solution are equation 2 and 3.

OpenStudy (anonymous):

thank you!! you helped a lot I have another one im gonna post if you could help me with it aswell? cx